OpenStudy (anonymous):

how to solve: lim (x-->5) (1/x-1/5)(2/x-5)

OpenStudy (anonymous):

I can't tell if you are supposed to have more brackets in there or not.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 5} (1/x - 1/5) (2/x-5)\]

OpenStudy (ash2326):

@azalea Why don't you just put x=5 and see what you get???

OpenStudy (jlastino):

First you take the limit of the function to see if it would have a value or not (1/5-1/5)(2/5-5) (0 . infinity) It is in an indeterminate form so combine the fractions (10-2x)/(5x^2-25x) Then take the limit 10-2(5)/5(5^2)-25(5) =0/0 (Still indeterminate) Use lhospitals rule take derivative of numerator / take derivative of denominator -2/10x-25 Now evaluate the lmit -2/10(5)-25 -2/50-25 -2/25

OpenStudy (anonymous):

@ash2326 because it gives me 0

OpenStudy (anonymous):

0 is a correct limit. The only problem is when you get an error (aka an indeterminate form).

OpenStudy (anonymous):

i have problems with knowing when it is indeterminate and when to use lhosiptals rule

OpenStudy (anonymous):

Ok, so you need to use L'Hopital's rule when you have the form: \[\frac {0}{0} \text{ or } \frac {\pm \infty}{0} \text{ or } \frac{\pm \infty} {\pm \infty}\]

OpenStudy (anonymous):

Those are all the indeterminate forms.

OpenStudy (anonymous):

thank you! I'll keep that in mind

OpenStudy (anonymous):

Also keep in mind that if you have: \[\infty \times 0\] You need to make it into the form: \[\infty \times 0 = \frac {\infty} {\frac{1}{0}} = \frac {\infty}{\infty}\] So you are then able to use L'Hopital's rule.

OpenStudy (anonymous):

oh ok! NOw i know why i can never solve some problems

OpenStudy (anonymous):

That should help a lot then :)

OpenStudy (anonymous):

i have another limits problem

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} x-\sqrt{x ^{2} +2x -4}\]

OpenStudy (anonymous):

First step: factor out an x.

OpenStudy (ash2326):

@azalea close this question and post a new question. That way you'll get more users to answer

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Then you use the step that I showed you above ;)